Circular No. 6651

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRS 1915+105 C. R. Robinson, S. N. Zhang, and M. L. McCollough, Universities

Space Research Association (USRA) and Marshall Space Flight Center (MSFC); B. A. Harmon, MSFC, NASA; S. Dieters and W. S. Paciesas, University of Alabama at Huntsville; M. Tavani, Columbia University and IFCTR, Milan; R. P. Fender, Sussex University; G. G. Pooley, Cambridge University; E. Waltman, Naval Research Laboratory; I. F. Mirabel, Service d’Astrophysique, Centre d’Etudes de Saclay; R. M. Hjellming, M. Rupen, and F. Ghigo, National Radio Astronomy Observatory; L. F. Rodriguez, Universidad Nacional Autonoma de México; and K. Ebisawa, USRA and Goddard Space Flight Center, report: “Following recent flaring activity at 2.25 and 8.3 GHz observed by the Green Bank Interferometer, and hard-x-ray activity observed by the Compton Gamma Ray Observatory (+ BATSE), we observed the superluminal jet source GRS 1915+105 during May 7.90–8.03 UT with the RXTE/PCA. Four repeating outburst cycles, of period near 2000 s, were observed from the source. The gross structure of the light curve can be described as a high-intensity spike, reaching a height of 3200 counts/s, followed by lower-intensity oscillations, a dip to a minimum level, and finally a gradual recovery after which the entire cycle repeats. The hardness ratio (defined above and below 11 keV) is inversely correlated with intensity. BATSE monitoring of GRS 1915+105 before and during these observations indicates a persistent hard component with intensity 0.10–0.13 photon cm^{−2} s^{−1} in the band 20–100 keV. Preliminary analysis of data since mid-April indicates that the energy spectrum can be adequately fit with a power law of photon index ∼ −3. The day-to-day variability of the source increases roughly in conjunction with the first GBI flare on April 25–26. Additional RXTE observations and a coordinated infrared/VLA/VLBA observation is scheduled for May 15.54–15.67. A light curve from part of our RXTE observations, future observing plans, and additional information on galactic superluminal sources will be available from

the WWW site at http://www.batse.msfc.nasa.gov/multiwave. Inter-

ested observers are encouraged to visit this site and to send information on

their observations to multiwave@bbking.msfc.nasa.gov.”

SUPERNOVA 1997bp IN NGC 4680 Unfiltered CCD magnitude estimate by S. Moretti and S. Tomaselli,

Forlí, Italy: May 2.94 UT, 15.2 (GSC comparison stars).

1997 May 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green